OpenStudy (anonymous):

Find the area of region bounded between the curves f(x)=cosx andg(x)=sinx on the interval [0,pi/2]

OpenStudy (anonymous):

\[\int\limits_{0}^{(\pi)/2}[cosx-sinx]dx\]

OpenStudy (anonymous):

do it

OpenStudy (anonymous):

that will give an answer of 0... the answer is 2 * int(cosx) from pi/4 to pi/2 = 2 * [sinx]pi/2,pi/4 = 2(1 - sqrt(2)/2) = 2 - sqrt(2)

OpenStudy (amistre64):

\[2\int_{0}^{pi/4}\color{red}{cos(x)}-\color{green}{sin(x)}\ dx\] \[\color{red}{sin(x)}+\color{green}{cos(x)}\] good eye james :) when sin and cosine migrate over each other, the area between them negates the area before them.

OpenStudy (anonymous):

question should be more specific...does it mean between the vertical boundaries of the two curves (amistre's solution) or between the horizontal boundaries of the two curves (my solution)?